# Random local access for sampling $k$-SAT solutions

### v1

@@ 2.1 The random local access model and local computation algorithms
One of the most widely studied models of local computation are local computation algorithms (LCAs) introduced in [1, 25]. Given a computation problem $F$, an LCA (in an online fashion) provides the $i$-th bit to some consistent solution $F$ in sublinear time. As originally defined, local computation algorithms need not be memory-less; in other words, the output solution can depend on the order of queried bits. However, several follow-up works have given query-oblivious or memory-less analogues of LCAs for a variety of natural problems. Such models are the non-random-sampling version of the random local access models we study here.
In this work, we construct a query-oblivious LCA for an intermediate step in our construction of a random local access model (described in more detail in Section 2.3). We thus precisely define both LCAs and random local access models below.
@@
Given an object family $(\Pi,\Omega)$ with input $\Pi$, sample space $\Omega\subseteq\Sigma^{n}$ (for some alphabet $\Sigma$), a $(t(n),\delta(n))$-local computation algorithm (LCA) provides an oracle $\mathcal{A}$ that implements query access to some arbitrary $\sigma\in\Omega$ that satisfies the following conditions:
@@
We say that $\mathcal{A}$ is a $(t(n),\delta(n))$-strongly local computation algorithm if $t(n)\lesssim\log^{c}(n)$ for some constant $c>0.$ We further say that $\mathcal{A}$ is query-oblivious if the outputs of $\mathcal{A}$ do not depend on the order of the queries but depend only on $\Pi$ and $\mathbf{R}$.
@@
Given a random object family $(\Pi,\Omega,\mu)$ with input $\Pi$, sample space $\Omega\subseteq\Sigma^{n}$ (with alphabet $\Sigma$), and distribution $\mu$, a $(t(n),\delta(n))$-random local access implementation of a family of query functions $\{F_{1},F_{2},\ldots\}$ provides an oracle $\mathcal{A}$ with the following properties:
@@
    $\|\mathsf{Law}(Y)-\mu\|_{\operatorname{\mathsf{TV}}}=\frac{1}{2}\sum_{\omega\in\Omega}\|\mathbb{P}(Y=\omega)-\mu(\omega)\|<\delta(n),$
@@ 2.2 Marking and shattering
  • While there exists a clause $C\in\mathcal{C}(\Phi)$ that is violated by the current assignment, resample $\operatorname{vbl}(C)$ uniformly at random.
The Lovász local lemma can be made quantitative, showing that not only is there some satisfying assignment to $\Phi$ if $d\leq(2ek)^{-1}2^{k}$, but both that there are exponentially many such satisfying assignments and that the marginal distributions $\mu_{v}$ are approximately uniform with $\mu_{v}(1)\leq\frac{1}{2}\exp(1/k)$ (see [21, 8]). Such local uniformity is critical to the success of algorithms that approximately sample from $\mu$. In their breakthrough work, Moitra [21] noted that this local uniformity continues to hold for conditional distributions $\mu^{\sigma}$ provided that each clause has a sufficiently large number of free variables under partial assignment $\sigma$. This motivates the idea of a marking, as introduced in [21], which is a careful selection (via the Lovász local lemma) of a linear sized subset of variables $\mathcal{M}\subseteq V$ that has the following properties:
  • For every clause $C$, $\mathcal{M}\cap\operatorname{vbl}(C)\gtrsim k$. Having a large number of marked variables in each clause would hopefully result in a desired shattering condition, namely, we can sample a partial assignment $\sigma:\mathcal{M}\to\{0,1\}$ on this marking that partitions the original formula into sufficiently small connected components.
  • For every clause $C$, $\operatorname{vbl}(C)\backslash\mathcal{M}\gtrsim k$. This large non-intersection preserves the local uniformity of the marginal distributions of the variables in $\mathcal{M}$.
The general strategy of arguments following the marking approach is to show that it is “easy” to sample a partial assignment $\sigma:\mathcal{M}\to\{0,1\}$, and moreover, conditioned on any such assignment, the dependency graph $G_{\Phi^{\sigma}}$ is very likely to shatter into sufficiently small connected components such that the remaining variable assignments can be efficiently sampled from the conditional distribution by solving a smaller instance of the original problem.
We make these notions precise and give a simple condition under which there exists a linear-time algorithm to construct such a marking.
@@
Lemma 2.9 ([8]).
Given $(k,d)$-formula $\Phi$ and constant $\alpha\in(0,1/2)$, if $d\lesssim\frac{1}{\mathrm{poly}(k)}2^{k\frac{(1-2\alpha)^{2}}{6\ln 2}}$, then there exists an $\alpha$-marking, and one such marking can be found via the Moser-Tardos algorithm.
We further note that the desired shattering occurs with high probability conditioned on a partial assignment on the marked variables that follows the marginal distribution $\mu_{\mathcal{M}}$.
Definition 2.10 (Shattering).
Given $(k,d)$-formula $\Phi$. We say that assignment $\sigma:\mathcal{M}\to\{0,1\}$ $\delta$-shatters $\Phi$ if for all $v\in V$, the connected component in $G_{\Phi}^{\sigma}$ that contains $v$ has size $\lesssim dk\log(n/\delta).$
Lemma 2.11 ([8]).
If $\Phi$ is a $(k,d)$-formula and $\mathcal{M}$ is a valid $\alpha$-marking for constant $\alpha\in(0,1/2)$, then for $2^{\alpha k}\gtrsim d^{3}k^{3}$, we have
$\mathbb{P}_{\tau\overset{\mathrm{\mathbf{R}}}{\leftarrow}\mu_{\mathcal{M}}}[\tau\,\delta\text{-shatters }\Phi]\geq 1-\frac{\delta}{n^{2}}.$
In this work, we locally, greedily construct an $\alpha$-marking $\mathcal{M}$ using a local implementation of Moser-Tardos. We will further argue that because of the shattering property, we can locally compute the connected component of $G_{\Phi}^{\sigma}$ for some $\sigma\sim\mu_{\mathcal{M}}$ that contains a given vertex, without having to actually assign linearly many vertices.
2.3 Query-oblivious LCA for marking
Per above, we construct a query-oblivious LCA, $\mathsf{IsMarked}(\cdot)$, where for a $(k,d)$-formula $\Phi$ and a given range of $\alpha\in(0,1)$, $\mathsf{IsMarked}(\cdot)$ can take in any variable $v\in V$ and output either 0 or 1 indicating whether $v$ is in some $\alpha$-marking of $V$. Moreover, $\mathsf{IsMarked}(\cdot)$ takes sublinear time and when queried on all of $V$, gives a consistent $\alpha$-marking of $V$.
@@ 2.3 Query-oblivious LCA for marking
$16k^{4}d^{5}\leq 2^{(\beta_{2}-\beta_{1})k-h(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}})(1-\beta_{1})k},$ (2)
$\delta\mapsto(\beta_{2}-\delta)-h(\frac{\beta_{2}-\delta}{1-\delta})(1-\delta)\text{ is decreasing on $0\leq\delta\leq\beta_{1}$,}$
$2e(kd+1)<2^{(1-h(\frac{\alpha}{1-\beta_{2}}))(1-\beta_{2})k}.$
(Here $h(\cdot)$ denotes the binary entropy function $x\mapsto-x\log_{2}(x)-(1-x)\log_{2}(1-x)$.)
Then there exists a $\mathrm{polylog}(n)$ time local computation algorithm $\mathsf{IsMarked}(\cdot)$ which, given any variable $v\in V$, returns an assignment $s_{v}\in\{0,1\}$ denoting whether $v$ is contained in an $\alpha$-marking of $\Phi$. Moreover, with probability at least $1-1/n$, the response for every $v\in V$ gives a consistent $\alpha$-marking of $\Phi$.
The construction of $\mathsf{IsMarked}(\cdot)$ and the verification that it is a query-oblivious LCA is inspired by the approach used in [1] to obtain an oblivious LCA for hypergraph $2$-coloring. Thus, we defer the proof of Theorem 2.12 to Appendix A.
3 A random local access algorithm for $k$-CNF solutions
In this section, we introduce an algorithm (Algorithm 1) that locally accesses the assignment of a variable $v\in V$ in a uniformly random solution $\sigma$ to $\Phi$. Overall, given $k$-CNF formula $\Phi=(V,\mathcal{C})$ and variable $v\in V$, the algorithm performs the following:
@@ 3 A random local access algorithm for $k$-CNF solutions
    Locally decides whether $v$ lies an $\alpha$-marking of $\Phi$ ($\mathsf{IsMarked}(v)$), such that the response for every $v\in V$ gives a consistent $\alpha$-marking $\mathcal{M}\subseteq V$.
@@
    If $v$ is marked, then we sample $\sigma(v)$ by computing $\mathsf{MarginSample}(\Phi,\emptyset,v)$ (adapted from [16]), which may make further recursive calls to sample other marked variables.
@@
    If $v$ is not marked, then we perform a depth-first search starting from $v$ in the reduced dependency hypergraph to compute the restriction of $\sigma$ on the marked variables. We start from $\sigma=\emptyset$; for every marked variable $w$ we encounter that we have not yet sampled, we compute $\mathsf{MarginSample}(\Phi,\sigma,w)$ and update $\sigma(w)$, to eventually obtain a connected component $\Phi_{v}^{\sigma}$ that contains $v$ (that we will prove is polylogarithmic in size with high probability). We then perform a $\mathsf{UniformSample}$ algorithm (Theorem 4.3) to sample a uniformly random satisfying assignment $\sigma$ for this connected component, and output $\sigma(v)$.
Algorithm 1 is our main routine; it has several subroutines which we describe in detail below.
Throughout the algorithm, we use public randomness $\mathbf{R}$ to fix a random global ordering of the variables $v_{1},\dots,v_{n}$ (e.g., by assigning an independent random value in $[0,1]$ to each $v_{i}$), such that one can locally query the relative ordering between any two variables in constant time.
@@
1:Compute $\mathsf{IsMarked}(v)$
2:if $\mathsf{IsMarked}(\Phi,v)$ then
3:     return $\mathsf{MarginSample}(\Phi,\emptyset,v)$
4:else
5:     $\Phi_{v}\leftarrow\mathsf{Conn}(\Phi,\emptyset,v)$
6:     $\sigma\leftarrow\mathsf{UniformSample}(\Phi_{v})$
7:     return $\sigma(v)$
We define (as in [16])
$\theta:=\frac{1}{2}-((1-2e\xi)^{-dk}-1)-(32ek^{2}d)^{-1},\quad\xi=\frac{1}{72e^{2}k^{3}d^{2}}.$ (3)
For a sufficiently sparse $(k,d)$-CNF $\Phi$, $\theta$ will be a pointwise lower bound on the marginal probabilities $\mu_{v}(\cdot)$. The basis of our local sampler will be the observation that $\theta\approx\frac{1}{2}$. Thus, we can sample the assignment of $v$ from $\mu_{v}$ via the following procedure: with probability $2\theta$, we uniformly at random select $v$ from $\{0,1\}$, and with probability $1-2\theta$, we sample from the overflow distribution $\mathcal{D}:=\frac{\mu_{v}^{\sigma}-\theta}{1-2\theta}$. Note that the first step can be completed in $O(1)$ time without attempting to sample any other variable spin. Thus, since $2\theta$ is relatively close to $1$, we can hope to locally sample $v$ without sampling too many other variable assignments. We make this idea more precise below with the $\mathsf{MarginSample}$ (Algorithm 2) and $\mathsf{MarginOverflow}$ (Algorithm 3) algorithms, adapted from [16].
Algorithm 2 $\mathsf{MarginSample}(\Phi,\sigma,v)$
Input: $k$-CNF formula $\Phi=(V,\mathcal{C})$, $v\in V$ with $\mathsf{IsMarked}(v)=1$, partial assignment $\sigma$
      Output: random assignment $\sigma(v)\in\{0,1\}$
@@
4:else
5:     $\sigma(v)\leftarrow\mathsf{MarginOverflow}(\Phi,\sigma_{v\leftarrow\star},v)$
6:return $\sigma(v)$
Algorithm 3 $\mathsf{MarginOverflow}(\Phi,\sigma,v)$
Input: $k$-CNF formula $\Phi=(V,\mathcal{C})$, $v\in V$ with $\mathsf{IsMarked}(v)=1$, partial assignment $\sigma$
      Output: random assignment $\sigma(v)\in\{0,1\}$
1:Draw $u\leftarrow\mathsf{NextVar}(\sigma)$ (3.1)
2:if $u\neq\perp$ then
@@
4:     if $r_{u}<2\theta$ then

### v2

@@ 2.1 The random local access model and local computation algorithms
One of the most widely studied models of local computation are local computation algorithms (LCAs) introduced in [1, 28]. Given a computation problem $F$, an LCA (in an online fashion) provides the $i$-th bit to some consistent solution $F$ in sublinear time. As originally defined, local computation algorithms need not be query-oblivious; in other words, the output solution can depend on the order of queried bits. However, several follow-up works have given query-oblivious analogues of LCAs for a variety of natural problems. Such models are the non-random-sampling version of the random local access models we study here.
In this work, we construct a query-oblivious LCA for an intermediate step in our construction of the random local access model (described in more detail in Section 2.2). We thus precisely define both LCAs and random local access models below.
@@
Given an object family $(\Pi,\Omega)$ with input $\Pi$ and sample space $\Omega\subseteq\Sigma^{n}$ (for some alphabet $\Sigma$), a $(t(n),\delta(n))$-local computation algorithm (LCA) provides an oracle $\mathcal{A}$ that implements query access to some arbitrary $\sigma\in\Omega$ that satisfies the following conditions:
@@
We further say that $\mathcal{A}$ is query-oblivious if the outputs of $\mathcal{A}$ do not depend on the order of the queries but depend only on $\Pi$ and $\mathbf{R}$.
@@
Given a random object family $(\Pi,\Omega,\mu)$ with input $\Pi$, sample space $\Omega\subseteq\Sigma^{n}$ (with alphabet $\Sigma$) and distribution $\mu$ supported on $\Omega$, a $(t(n),\delta(n))$-random local access implementation of a family of query functions $\{F_{1},F_{2},\ldots\}$ provides an oracle $\mathcal{A}$ with the following properties:
@@
    $d_{\operatorname{\mathsf{TV}}}(Y,\mu)=\frac{1}{2}\sum_{\omega\in\Omega}|\mathbb{P}(Y=\omega)-\mu(\omega)|<\delta(n),$
@@ 2.2 Marking and a query-oblivious LCA
  • While there exists a clause $C\in\mathcal{C}$ that is violated by the current assignment, resample variables in $C$ uniformly at random.
The Lovász local lemma can be made quantitative, showing that not only is there some satisfying assignment to $\Phi$ if $d\leq(2ek)^{-1}2^{k}$, but both that there are exponentially many such satisfying assignments and that the marginal distributions $\mu_{v}$ are approximately uniform with $\mu_{v}(1)\leq\frac{1}{2}\exp(1/k)$ (see [24, 10]). Such local uniformity is critical to the success of algorithms that approximately sample from $\mu$. In his breakthrough work, Moitra [24] noted that this local uniformity continues to hold for conditional distributions $\mu^{\sigma}$ provided that each clause has a sufficiently large number of free variables under partial assignment $\sigma$. This motivates the idea of a marking, as introduced in [24], which is a careful selection (via the Lovász local lemma) of a linear sized subset of variables $\mathcal{M}\subseteq V$ that has the following properties:
  • For every clause $C$, $|\operatorname{vbl}(C)\cap\mathcal{M}|\gtrsim k$. Having a large number of marked variables in each clause would hopefully result in a desired shattering condition; namely, we can sample a partial assignment $\sigma\in\{0,1\}^{\mathcal{M}}$ on this marking that partitions the original formula into sufficiently small connected components.
  • For every clause $C$, $|\operatorname{vbl}(C)\backslash\mathcal{M}|\gtrsim k$. This large non-intersection preserves the local uniformity of the marginal distributions of the as yet-unsampled variables in $\mathcal{M}$.
The general strategy of arguments following the marking approach is to show that it is “easy” to sample a partial assignment $\sigma\in\{0,1\}^{\mathcal{M}}$, and moreover, conditioned on any such assignment, the reduced formula $\Phi^{\sigma}$ is very likely to shatter into sufficiently small connected components such that the remaining variable assignments can be efficiently sampled from the conditional distribution by solving a smaller instance of the original problem. We now make this notion precise.
@@
In this work, we locally, greedily construct an $\alpha$-marking $\mathcal{M}$ using a local implementation of Moser-Tardos. We will further argue that because of the shattering property, we can locally compute the connected component of $\Phi_{v}^{\sigma}$ for some $\sigma\sim\mu_{\mathcal{M}}$ and a given vertex $v$, without having to actually assign linearly many vertices.
Precisely, we construct a query-oblivious LCA, $\mathsf{IsMarked}(\cdot)$, where for a $(k,d)$-formula $\Phi$ and a given range of $\alpha\in(0,1)$, $\mathsf{IsMarked}(\cdot)$ can take in any variable $v\in V$ and output either 0 or 1 indicating whether $v$ is in some $\alpha$-marking of $V$. Moreover, $\mathsf{IsMarked}(\cdot)$ takes sublinear time and when queried on all of $V$, gives a consistent $\alpha$-marking of $V$.
@@
$16k^{4}d^{5}\leq 2^{(\beta_{2}-\beta_{1})k-h\left(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}}\right)(1-\beta_{1})k},$ (2)
$\delta\mapsto(\beta_{2}-\delta)-h\left(\frac{\beta_{2}-\delta}{1-\delta}\right)(1-\delta)\text{ is decreasing on $0\leq\delta\leq\beta_{1}$,}$
$2e(kd+1)<2^{\left(1-h\left(\frac{\alpha}{1-\beta_{2}}\right)\right)(1-\beta_{2})k}.$
(Here $h(x):=-x\log_{2}(x)-(1-x)\log_{2}(1-x)$ is the binary entropy.)
Fix constant $c>0$. Then there exists a $\mathrm{polylog}(n)$ time local computation algorithm $\mathsf{IsMarked}(\cdot)$ which, given any variable $v\in V$, returns an assignment $s_{v}\in\{0,1\}$ denoting whether $v$ is contained in an $\alpha$-marking of $\Phi$. Moreover, with probability at least $1-n^{-c}$, the responses for all $v\in V$ yield a consistent $\alpha$-marking of $\Phi$.
The construction of $\mathsf{IsMarked}(\cdot)$ and the verification that it is a query-oblivious LCA draws inspiration from the approach in [1] to obtain an oblivious LCA for hypergraph $2$-coloring. We defer the proof of Theorem 2.10 to Appendix D.
3 A random local access algorithm for $k$-SAT solutions
In this section, we introduce the main algorithm (Algorithm 1) that locally accesses the assignment of a variable $v\in V$ in a uniformly random satisfying assignment $\sigma$ to $\Phi$.
Recall from 1.1 that given a $(k,d)$-formula $\Phi=(V,\mathcal{C})$, variable $v\in V$, and any constant $c>0$, we wish to come up with a random local access algorithm $\mathcal{A}$ such that (1) the expected running time of $\mathcal{A}$ is $\mathrm{poly}\log(n)$, and (2) the output $\widehat{\mu}_{v}$ of $\mathcal{A}$ for every $v\in V$ satisfies $d_{\operatorname{\mathsf{TV}}}(\textsf{Law}(\widehat{\mu}_{v}),\mu_{v})\leq\frac{1}{n^{c}}$. As a high level description, given the above inputs, Algorithm 1 performs the following:
@@ 3 A random local access algorithm for $k$-SAT solutions
    Locally decides whether $v$ lies in an $\alpha$-marking of $\Phi$ using $\mathsf{IsMarked}(v)$ (Theorem 2.10), such that the responses for all $v\in V$ yield a consistent $\alpha$-marking $\mathcal{M}\subseteq V$.
@@
    Suppose $\sigma\sim\mu$ is a uniformly random satisfying assignment to $\Phi$. If $v$ is marked, then we sample $\sigma(v)$ by computing $\mathsf{MarginSample}(v)$ (adapted from [9, Algorithm 8]), which may make further recursive calls to sample other marked variables.
@@
    If $v$ is not marked, then we perform a depth-first search starting from $v$ to compute $\sigma$ restricted to $\mathcal{M}$. We start from $\sigma=\emptyset$; for every $w\in\mathcal{M}$ we encounter that we have not yet sampled, we compute $\mathsf{MarginSample}(w)$ and update $\sigma(w)$, to eventually obtain a (w.h.p. polylogarithmic in size) connected component $\Phi_{v}:=\Phi_{v}^{\sigma}$ that contains $v$. This part is carried out by the algorithm $\mathsf{Conn}(v)$.
    After obtaining the connected component $\Phi_{v}^{\sigma}$, we call $\mathsf{UniformSample}$ (Theorem 3.1) to sample a uniformly random satisfying assignment to $\Phi_{v}^{\sigma}$ and extend $\sigma$. We then output $\sigma(v)$.
@@
1:if $\mathsf{IsMarked}(v)$ then
2:     return $\mathsf{MarginSample}(v)$
3:else
4:     $\Phi_{v}\leftarrow\mathsf{Conn}(v)$
5:     $\sigma\leftarrow\mathsf{UniformSample}(\Phi_{v})$
6:     return $\sigma(v)$
Algorithm 1 is our main routine; it has several subroutines $\mathsf{IsMarked}$, $\mathsf{MarginSample}$, $\mathsf{Conn}$, and $\mathsf{UniformSample}$ that we now describe. Recall that $\mathsf{IsMarked}$ has been discussed in Theorem 2.10. We now introduce the algorithm $\mathsf{UniformSample}$ given by the work of He–Wang–Yin [18].
Theorem 3.1 ([18, Theorems 5.1 and 6.1]).
Suppose $\Phi$ is a $(k,d)$-CNF formula on $n$ variables with $k\cdot 2^{-k}\cdot(dk)^{5}\cdot 4\leq\frac{1}{150e^{3}}$. Then there exists an algorithm $\mathsf{UniformSample}$ that terminates in $O(k^{3}(dk)^{9}n)$ time in expectation, and outputs a uniformly random satisfying assignment to $\Phi$.
As seen in Theorem 3.1, for an $n$-variable $(k,d)$-CNF formula, the algorithm has running time $O(n)$. However, as we will only apply $\mathsf{UniformSample}$ to connected components of size $\mathrm{poly}\log(n)$ (that are shattered by the partial assignment on the marked variables), the execution of $\mathsf{UniformSample}$ in Line 6 of Algorithm 1 will only takes polylogarithmic time.
We will define subroutines $\mathsf{MarginSample}$ and $\mathsf{Conn}$ below and show that they satisfy the desired correctness and efficiency properties, beginning by verifying $\mathsf{MarginSample}$ has the desired properties in Section 4.
Theorem 3.2.
Let $\Phi=(V,\mathcal{C})$ be a $(k,d)$-formula, $\alpha>0$, and $\mathcal{M}\subseteq V$ be an $\alpha$-marking as in 2.9. Fix positive constant $c>0$. Suppose $\theta:=1-\frac{1}{2}\exp\left(\frac{2edk}{2^{\alpha k}}\right)\geq 0.4$, $236ed^{3}k^{4}\cdot 0.6^{\alpha k}\leq 1/$ 2, and $2^{-\frac{1}{48dk^{4}}}\cdot e^{\frac{2d^{2}/\alpha}{2^{\alpha k}}}\leq 0.9$. Then there exists a random local access algorithm $\mathsf{MarginSample}(\cdot)$ such that for every $u\in\mathcal{M}$, $\mathsf{MarginSample}(u)$ returns a random value in $\{0,1\}$ that satisfies the following:
  1. 1.
    Let $\nu:=\mu_{\mathcal{M}}$ and $\widehat{\nu}$ be the joint distribution of the outputs $(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}$. Then we have $d_{\operatorname{\mathsf{TV}}}(\widehat{\nu},\nu)<n^{-c}$.
  2. 2.
    For every $u\in\mathcal{M}$, the expected cost of $\mathsf{MarginSample}(u)$ is $\mathrm{poly}\log(n)$.
We also require $\mathsf{Conn}$ to be correct and have low expected cost, verified in Appendix C.
Theorem 3.3.
Let $\Phi=(V,\mathcal{C})$ be a $(k,d)$-formula, $\alpha>0$, and $\mathcal{M}\subseteq V$ be an $\alpha$-marking as in 2.9. Suppose $\theta:=1-\frac{1}{2}\exp\left(\frac{2edk}{2^{\alpha k}}\right)\geq 0.4$ and $d\leq 2^{\alpha k/4}$. Then there exists a random local access algorithm $\mathsf{Conn}(\cdot)$ such that for every $v\in V\setminus\mathcal{M}$, $\mathsf{Conn}(v)$ returns the connected component in $\Phi^{\sigma}$ that contains $v$, where $\sigma$ the partial assignment given by $(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}$. Moreover, for every $v\in V\setminus\mathcal{M}$, the expected cost of $\mathsf{Conn}(v)$ is $\mathrm{poly}\log(n)$.
4 Proof of Theorem 3.2
In this section we show Theorem 3.2. Throughout, let $\Phi=(V,\mathcal{C})$ be a $(k,d)$-formula, $\alpha>0$, and $\mathcal{M}\subseteq V$ be an $\alpha$-marking with $|\mathcal{M}|=m$. We introduce a local access algorithm $\mathsf{MarginSample}$ that satisfies the key property that the joint distribution of outputs $(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}$ consistently follows the marginal distribution $\mu_{\mathcal{M}}$.
In particular, $(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}$ will simulate the output of a systematic scan Glauber dynamics on the marked variables. We begin by recalling some basics of Markov chains and the projected Glauber dynamics Markov chain for sampling $k$-SAT solutions, introduced in [10].
Definition 4.1.
A Markov chain $(X_{t})_{t\geq 0}$ over a state space $\Omega$ is given by a transition matrix $P:\Omega\times\Omega\to\mathbb{R}_{\geq 0}$. A distribution $\mu$ over $\Omega$ is a stationary distribution of $P$ if $\mu=\mu P$. The Markov chain $P$ is reversible with respect to $\mu$ if it satisfies the detailed balance condition $\mu(X)P(X,Y)=\mu(Y)P(Y,X)$, which implies that $\mu$ is a stationary distribution for $P$.
We say that a Markov chain $P$ is irreducible if for any $X,Y\in\Omega$, there exists some integer $t$ so that $P^{t}(X,Y)>0$ and aperiodic if $\gcd\{t\mid P^{t}(X,X)>0\}=1$. If $P$ is irreducible and aperiodic, then it converges to unique stationary distribution $\mu$ with mixing time
$T_{\text{mix}}(P,\varepsilon)=\max_{X_{0}\in\Omega}\min\left\{t:d_{\operatorname{\mathsf{TV}}}(P^{t}(X_{0},\cdot),\mu)\leq\varepsilon\right\}.$
We now describe one specific Markov chain, the projected Glauber dynamics on the marked variables, as studied by Feng–Guo–Yin–Zhang [10]. They gave the first Markov chain-based algorithm to sample from the uniform distribution of satisfying assignments to a given bounded degree $k$-CNF at exponential clause density, following earlier work of Moitra [24] who gave the first fixed-parameter tractable algorithm to sample from this distribution. Since these works, there has been a tremendous amount of progress on sampling from atomic CSPs in the local lemma regime [2, 6, 9, 13, 12].
Algorithm 2 Projected Glauber dynamics
Input: a $k$-CNF formula $\Phi=(V,\mathcal{C})$, a set of marked variables $\mathcal{M}\subseteq V$, and time $T$
      Output: a random assignment $X_{*}\in\{0,1\}^{\mathcal{M}}$
@@
4:     For all $u^{\prime}\in\mathcal{M}\setminus\{u\}$, $X_{t}(u^{\prime})\leftarrow X_{t-1}(u^{\prime})$
@@
6:Return $X_{*}=X_{T}$
We contrast Algorithm 2 with the systematic scan version of the projected Glauber dynamics. Instead of resampling a random variable at time $t$, in the systematic scan Glauber dynamics, we always resample the variable $u_{i(t)}$ at time $t$, where $i(t)$ is a deterministic function of $t$.
Definition 4.2.
Let =$\mathcal{M}=\{u_{1},\ldots,u_{m}\}$ denote the marked variables (so $|\mathcal{M}|=m$).
  • For every time $t\in\mathbb{Z}$, define $i(t):=(t\mod m)+1$.
  • For every time $t\in\mathbb{Z}$ and $u_{i}\in\mathcal{M}$, define $\mathsf{pred}_{u_{i}}(t):=\max\{s\leq t:i(s)=i\}.$
In the systematic scan Glauber dynamics, we always resample vertex $u_{i(t)}$ at time $t$ (as opposed to randomly choosing a variable to resample at each step). For every $u\in\mathcal{M}$ and time $t$, $\mathsf{pred}_{u}(t)$ denotes the most recent time up to time $t$ at which $u$ got resampled. Observe that for all $t\in\mathbb{Z}$ and $u\in\mathcal{M}$, we have $t-m<\mathsf{pred}_{u}(t)\leq t$. Moreover, for all $w\neq u_{i(t)}$, we have $\mathsf{pred}_{w}(t)<t$.
Algorithm 3 Systematic scan projected Glauber dynamics
Input: a $k$-CNF formula $\Phi=(V,\mathcal{C})$, a set of marked variables $\mathcal{M}\subseteq V$, time $T$, and an ordering =$\mathcal{M}=\{u_{1},\ldots,u_{m}\}$
      Output: a random assignment $X_{*}\in\{0,1\}^{\mathcal{M}}$